# Period collapse of Markov triangles

### v1

@@ 1.1 Integral affine geometry
Here we define the notions of planar integral affine geometry used in this paper. We refer the interested reader to [10] for an extensive account of the theory. By a polygon $P$, we mean a two-dimensional convex polytope, i.e. the convex hull of finitely many points in $\mathbb{R}^{2}$, called the vertices of $P$, while the boundary line segments connecting vertices are called the edges of $P$. A polygon is said to be rational resp. integral if its vertices have rational resp. integral coordinates. For rational polygons we can define some additional notions: consider three distinct points $p_{1},p_{2},p_{3}\in\mathbb{R}^{2}$ with rational coordinates.
@@
Figure 1: Integrally congruent triangles.
@@ 1.2 Ehrhart theory
Figure 3: The limiting Fibonacci triangle.
@@ 1.3 Markov triangles
(1,1,1)(1,1,1)$(1,2,1)$ $(1,5,2)$ $(1,13,5)$ $(1,34,13)$ $(5,194,13)$ $(2,29,5)$ $(5,433,29)$ $(2,169,29)$
Figure 4: The first five generations of the Markov tree.
Let $p\notin\{1,2\}$ be a Markov number and $\mathfrak{m}$ a Markov triple in which $p$ is the largest number. We denote by $\Lambda(\mathfrak{m})$ the subgraph of the Markov tree consisting of all triples obtained by an iteration of mutations of $\mathfrak{m}$ preserving $p$. $\Lambda(\mathfrak{m})$ is a bivalent subtree rooted in $\mathfrak{m}=\vcentcolon(a,p,b)$ consisting of two branches: the $B_{p}^{a}$-branch obtained by mutating $(a,p,b)$ first at $b$, and the $B_{p}^{b}$-branch obtained by mutating $(a,p,b)$ first at $a$, each of which then extends by mutations that preserve $p$ and the largest element of the triple. As an example, the subtree formed by the coloured vertices in Figure 4 is $\Lambda(1,5,2)$. The first two vertices of the branch $B_{5}^{1}$ are the red triples and those of the branch $B_{5}^{2}$ are the blue triples.
@@
Figure 5: Geometric mutation of a $(1,5,2)$-triangle at the $1$-vertex.
@@ 1.4 Ehrhart theory of Markov triangles
    Figure 6: A $(2,5,29)$-triangle in standard $2$-position.
@@
    Figure 7: $(2,5,29)$-triangles in a generic barycentric position (left) and in a standard $2$-barycentric position (right).
@@
Figure 8: Exaggerated representation of a sequence of Markov triangles in standard $a$-position.
@@
Recall from Example 1.9 the triangle $\mathcal{F}_{(b/c,c/b)}$ for integers $b,c\in\mathbb{N}$ with one vertex at the origin and the two other vertices at $\frac{b}{c}(0,1)$ resp. $\frac{c}{b}(1,0)$. Gardiner and Kleinman show in [5, Theorem 1.6] that $\mathcal{F}_{(b/c,c/b)}$ is pseudo-integral if and only if $b$ and $c$ are consecutive odd-indexed Fibonacci numbers. This motivates the following problem: take coprime integers $a,q\in\mathbb{N}$. For integers $b,c\in\mathbb{N}$, define the triangle $\mathcal{F}^{(p,q)}_{(b/c,c/b)}$ with one vertex at the origin and the two other vertices at $\frac{b}{c}(aq-1,a^{2})$ resp. $\frac{c}{b}(1,0)$. Proposition 1.21 shows that $\mathcal{F}^{(p,q)}_{(b/c,c/b)}$ is pseudo-integral if $(a,b,c)$ is a Markov triple and $q\equiv 3cb^{-1}$ mod $a$, since in this case $\mathcal{F}^{(p,q)}_{(b/c,c/b)}$ is the $a$-dilate of a Markov triangle representing $(a,b,c)$ in standard $a$-position. This generalizes one direction of [5, Theorem 1.6]. Is the converse true, i.e. do there exist pseudo-integral triangles $\mathcal{F}^{(p,q)}_{(b/c,c/b)}$ which are not Markov? A similar question may be asked in the barycentric setting: Proposition 1.20 says that the $\frac{3}{a}$-dilate of the $(\frac{-q}{3a},\frac{-1}{3})$-translate of the triangle $\mathcal{F}^{(p,q)}_{(b/c,c/b)}$ is pseudo-integral if $(a,b,c)$ is a Markov triple and $q\equiv 3cb^{-1}$ mod $a$. Is the converse true?
@@ 2.2 On model Markov triangles
Figure 10: Preservation of the integral barycentre under geometric mutation.
@@ 2.3 On the Ehrhart theory of rational Markov triangles
Figure 11: Geometric mutation of $\Delta$ at $m_{n}$.

### v2

@@ Abstract
Keywords.
Ehrhart theory, Markov triangles, period collapse, affine integral geometry, symplectic geometry.
Mathematics Subject Classification.
52C05 (53D35)
Acknowledgements.
@@ 1.1 Integral affine geometry
Here we define the notions of planar integral affine geometry used in this paper. We refer the interested reader to [10] for an extensive account of the theory. By a polygon $P$, we mean a two-dimensional convex polytope, i.e. the convex hull of finitely many points in $\mathbb{R}^{2}$, called the vertices of $P$, while the boundary line segments connecting vertices are called the edges of $P$. A polygon is said to be rational resp. integral if its vertices have rational resp. integral coordinates.
For rational polygons we can define some additional notions: consider three distinct points $p_{1},p_{2},p_{3}\in\mathbb{R}^{2}$ with rational coordinates.
@@
Figure 1: Examples of integrally congruent triangles
@@ 1.2 Ehrhart theory
Figure 3: The limiting Fibonacci triangle
@@ 1.3 Markov triangles
(1,1,1)(1,1,1)$(1,2,1)$ $\underline{\overline{(1,5,2)}}$ $\overline{(1,13,5)}$ $(1,34,13)$ $\overline{(5,194,13)}$ $\underline{(2,29,5)}$ $\underline{(5,433,29)}$ $(2,169,29)$
Figure 4: The first five generations of the Markov tree
Let $p\notin\{1,2\}$ be a Markov number and $\mathfrak{m}$ a Markov triple in which $p$ is the largest number. We denote by $\Lambda(\mathfrak{m})$ the subgraph of the Markov tree consisting of all triples obtained by an iteration of mutations of $\mathfrak{m}$ preserving $p$. $\Lambda(\mathfrak{m})$ is a bivalent subtree rooted in $\mathfrak{m}=\vcentcolon(a,p,b)$ consisting of two branches: the $B_{p}^{a}$-branch obtained by mutating $(a,p,b)$ first at $b$, and the $B_{p}^{b}$-branch obtained by mutating $(a,p,b)$ first at $a$, each of which then extends by mutations that preserve $p$ and the largest element of the triple. As an example, the subtree formed by the vertices which are either underlined or overlined in Figure 4 is $\Lambda(1,5,2)$. The first two vertices of the branch $B_{5}^{1}$ are the overlined triples and those of the branch $B_{5}^{2}$ are the underlined triples.
@@
Figure 5: Geometric mutation of a $(1,5,2)$-triangle at the $1$-vertex
@@ 1.4 Ehrhart theory of Markov triangles
    Figure 6: A $(2,5,29)$-triangle in standard $2$-position
@@
    Figure 7: $(2,5,29)$-triangles in a generic barycentric position (left) and in a standard $2$-barycentric position (right)
@@
Figure 8: Exaggerated representation of a sequence of Markov triangles in standard $a$-position
@@
Recall from Example 1.9 the triangle $\mathcal{F}_{(b/c,c/b)}$ for integers $b,c\in\mathbb{N}$ with one vertex at the origin and the two other vertices at $\frac{b}{c}(0,1)$ resp. $\frac{c}{b}(1,0)$. Cristofaro-Gardiner and Kleinman show in [5, Theorem 1.6] that $\mathcal{F}_{(b/c,c/b)}$ is pseudo-integral if and only if $b$ and $c$ are consecutive odd-indexed Fibonacci numbers. This motivates the following problem: take coprime integers $a,q\in\mathbb{N}$. For integers $b,c\in\mathbb{N}$, define the triangle $\mathcal{F}^{(p,q)}_{(b/c,c/b)}$ with one vertex at the origin and the two other vertices at $\frac{b}{c}(aq-1,a^{2})$ resp. $\frac{c}{b}(1,0)$. Proposition 1.21 shows that $\mathcal{F}^{(p,q)}_{(b/c,c/b)}$ is pseudo-integral if $(a,b,c)$ is a Markov triple and $q\equiv 3cb^{-1}$ mod $a$, since in this case $\mathcal{F}^{(p,q)}_{(b/c,c/b)}$ is the $a$-dilate of a Markov triangle representing $(a,b,c)$ in standard $a$-position. This generalizes one direction of [5, Theorem 1.6]. Is the converse true, i.e. do there exist pseudo-integral triangles $\mathcal{F}^{(p,q)}_{(b/c,c/b)}$ which are not Markov? A similar question may be asked in the barycentric setting: Proposition 1.20 says that the $\frac{3}{a}$-dilate of the $(\frac{-q}{3a},\frac{-1}{3})$-translate of the triangle $\mathcal{F}^{(p,q)}_{(b/c,c/b)}$ is pseudo-integral if $(a,b,c)$ is a Markov triple and $q\equiv 3cb^{-1}$ mod $a$. Is the converse true?
@@ 2.2 On model Markov triangles
Figure 10: Preservation of the integral barycentre under geometric mutation
@@ 2.3 On the Ehrhart theory of rational Markov triangles
Figure 11: Geometric mutation of $\Delta$ at $m_{n}$
@@ 2.4 On the Ehrhart theory of irrational Markov triangles
Figure 12: Representation of a neighbourhood of the symmetric difference $\Delta_{n}^{a}\ominus\Delta_{\infty}^{a}$